I have worked on this all day and seem not to understand where to put what to where, can you please help me?
the lumber company can convert logs into lumber or plywood, in a given day, the mill turns out twice as much units of plywood as lumber. It makes a profit of $20 on a unit of lumber and $40 on plywood. How many of each unit must be produced and sold in order to make a profit of $12520.

Answer by rfer(16322)

20x+40*2x=12520
100x=12520
x=125.2 lumber
2x=250.4 plywood
--------------------------
125.2*20=$2504
250.4*40=$10016